OpenStudy (anonymous):

Can someone help with with this, 15x-20y=7 5y=-3+10x

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Did you want to solve for x and y then?

OpenStudy (anonymous):

Assuming that you do, you want to multiply the bottom equation by 4, getting 20y = -12 + 40x, then you subtract 15x on both sides of the first one, getting -20y = 7-15x: After that you add both equations and solve for x, then plug x back in and solve for y x=1/11 y=-31/110